# You Only Sample Once: Taming One-Step Text-to-Image Synthesis by Self-Cooperative Diffusion GANs

### v4

@@ Appendix E Additional Related Work
Text-to-image Diffusion Models. Since Diffusion models have shown a stable training process and are well-suited for scaling up generative models, numerous works have been proposed to extend DMs for text-to-image generation [39, 1, 40, 61]. Latent DMs [40, 37] are widely adopted for high-resolution image and video generation due to their computational efficiency.
@@ G.1 Unconditional Generation Experiments
For the generator, we use the Adam optimizer with = 0.9 and = 0.999; for the discriminator, we use the Adam optimizer with = 0. and = 0.999. We adopt a constant learning rate of 2e-4 for both the discriminator and the generator. We apply EMA with a coefficient of 0.9999 for the generator. We let the and .

### v5

@@ Appendix E Additional Related Work
Text-to-image Diffusion Models. Since Diffusion models have shown a stable training process and are well-suited for scaling up generative models, numerous works have been proposed to extend DMs for text-to-image generation (Ramesh et al., 2022; Balaji et al., 2022; Rombach et al., 2022; Xue et al., 2023). Latent DMs (Rombach et al., 2022; Podell et al., 2023) are widely adopted for high-resolution image and video generation due to their computational efficiency.
@@ G.1 Unconditional Generation Experiments
For the generator, we use the Adam optimizer with = 0.9 and = 0.999; for the discriminator, we use the Adam optimizer with = 0. and = 0.999. We adopt a constant learning rate of 2e-4 for both the discriminator and the generator in training from scratch. We adopt a constant learning rate of 2e-5 for both the discriminator and the generator in fine-tuning. We apply EMA with a coefficient of 0.9999 for the generator. We let the and .
@@ G.2 Text-To-Image Experiments
For full fine-tuning, we train YOSO on the JourneyDB dataset (Pan et al., 2023), by resizing to 512 resolution. And we only use the square image. For LoRA fine-tuning, we use an internally collected dataset and the caption of the JourneyDB dataset (Pan et al., 2023) to generate data for training, we only generate one image for one caption. For the evaluation, we evaluate the HPS score on its benchmark, and we evaluate other metrics based on COCO-5k (Lin et al., 2014) datasets.
G.3 Ablation Study
For the ablation study on CIFAR-10, we use the same UNet and discriminator architecture as used in DDGANs (Xiao et al., 2022). For the generator, we use the Adam optimizer with = 0.9 and = 0.999; for the discriminator, we use the Adam optimizer with = 0. and = 0.999. We adopt a constant learning rate of 2e-4 for both the discriminator and the generator. We apply EMA with a coefficient of 0.9999 for the generator. We let the and .